Monday
This presentation is the property of its rightful owner.
Sponsored Links
1 / 11

Monday PowerPoint PPT Presentation


  • 89 Views
  • Uploaded on
  • Presentation posted in: General

Monday . Recover after trip of RB.A81  Cause of the trip not clear. MP3? 09:00-13:00 Re-inject and Injection verification. In parallel investigations on the injection pulse inhibit occurred on Sunday

Download Presentation

Monday

An Image/Link below is provided (as is) to download presentation

Download Policy: Content on the Website is provided to you AS IS for your information and personal use and may not be sold / licensed / shared on other websites without getting consent from its author.While downloading, if for some reason you are not able to download a presentation, the publisher may have deleted the file from their server.


- - - - - - - - - - - - - - - - - - - - - - - - - - E N D - - - - - - - - - - - - - - - - - - - - - - - - - -

Presentation Transcript


Monday

Monday

  • Recover after trip of RB.A81  Cause of the trip not clear. MP3?

  • 09:00-13:00 Re-inject and Injection verification. In parallel investigations on the injection pulse inhibit occurred on Sunday

  • 13:00-15:00 Pre-cycle of S12 after power cycle of nQPS crate B10L2 needs to be power cycled to re-establish local communication.

  • 15:00-17:00 Work in the injectors to adjust intensity, transverse, longitudinal parameters. RE-optimization of the scraping in the SPS to reduce losses (20% of beam lost)

  • 17:00 Beams dumped during injection of last but one 24 bunch train of B1 (losses on TCTH.4L2.B1)

  • 17:00-19:00 Reduction of the transverse blow-up in PS and SPS

  • 21:20 Stable beams #1408  Finally some physics!!

  • 07:00 End of physics #1408 (~2.4-2.5 pb-1 in 9.6 h) for vernier scans

  • 07:30 Beam dump

  • 08:30 Access


Injection

Injection

  • TI8 resteered in H and V to optimise losses and injection oscillations - corrected back to the reference and the injection looked good - IQC all green for 24b, oscillations below 0.5 mm

  • TI2 looks fine in steering - the -1.25 mm bump is left in through the MSI. Large oscillations in V as a result.

  • Checked effects of moving in some TI2 screens, on losses in MSI, with TCDIs out

  • Confirm the request to check locally the alignment of the MSI, especially in the region of transition between MSIB and MSIA.

B. Goddard, M. Meddahi


Injection kicker pulse inhibit

Injection kicker pulse inhibit

  • Drift of LHC timing wrt SPS timing due to GPS time reference.

  • LHC RF prepulse generation is based on SPS timing, but MKI resonant charging start is based on LHC timing.

  • Timing drift introduced a delay of roughly 200us on LHC injection forewarning (-1000ms) wrt RF injection prepulse. Problem disappeared after re-synchronization of the machines on Monday at 3 AM

Sat 09/10

Sun 10/10

Ph. Baudrenghien, E. Carlier, J. Kozsar, N. Magnin


Injection kicker pulse inhibit1

Injection kicker pulse inhibit

  • Actions:

    • Verification by expert of the status of the synchronization before injection (done for the past fill)

    • Replacement of the GPS timing receiver  this morning

    • Other actions under study:

      • pulse (instead of the LHC timing) to start charging the PFN, making the scheme more robust  we should go for that

      • extraction inhibit when injection pulse is inhibited


Stable beams 1408

Stable beams #1408

  • Initial emittances from luminosity ~2.5 um


Vacuum stable beams 1408

Vacuum Stable beams #1408

  • Switched OFF/ON solenoids at ~23:00?


Quasi ufo event

Quasi-UFO event

  • In 27.R7 ~10 mins after the start of physics21:33:31 - Warning on: BLMQI.27R7.B2I30_MQ, integration time: 640 us, losses = 9.250231E-02, threshold = 2.192476E-01, ratio = 42%21:33:31 - Warning on: BLMQI.27R7.B2I30_MQ, integration time: 2560 us, losses = 2.536262E-02, threshold = 5.481189E-02, ratio = 46%21:33:31 - Warning on: BLMQI.27R7.B2I20_MQ, integration time: 640 us, losses = 9.020587E-02, threshold = 2.192476E-01, ratio = 41%21:33:31 - Warning on: BLMQI.27R7.B2I20_MQ, integration time: 2560 us, losses = 2.487477E-02, threshold = 5.481189E-02, ratio = 45%21:33:31 - Warning on: BLMQI.27R7.B2I10_MQ, integration time: 320 us, losses = 3.716156E-01, threshold = 6.445071E-01, ratio = 58%21:33:31 - Warning on: BLMQI.27R7.B2I10_MQ, integration time: 640 us, losses = 2.699785E-01, threshold = 3.222535E-01, ratio = 84%21:33:31 - Warning on: BLMQI.27R7.B2I10_MQ, integration time: 2560 us, losses = 7.329793E-02, threshold = 8.056338E-02, ratio = 91%THAT WAS CLOSE!


Today

Today

  • Access from 08:30 to 12:00-13:00:

    • Access in IP2 for Injection area investigations (radiation survey, alignment, X-ray)

    • UJ33 for the installation of a magnetic probe at 09:00

    • Cryogenics in PM85 (no RP, no MAD required)

    • UPS in UX45

    • In parallel: change of GPS repeater

  • PM:

    • Recovery and injection verification (B1)

    • Abort gap cleaning at injection (4 hours)

  • Night: Physics fill with 248 bunches

  • Wednesday AM: VdM scan


Outstanding requests for studies

Outstanding requests for studies

  • Periodic loss maps (not really MD)

    • Betatron loss map done week 40.

  • Orbit shift versus TCT check: collide 1 bunch and check the TCT center change as compared to predicted orbit change.

  • BI

    • PLL studies during ramp

    • Commission BGI in preparation to ions – switch on at injection

    • Other BI studies – list to be updated

    • Test of 'dynamic' references for the OFB (varying Xing.separation bump): injection, then test of ramp + first part of squeeze.

  • BT

    • Abort Gap Cleaning – on at injection for physics

    • Injection Studies 32 Bunches when required (344 bunches)

  • Setting-up of the SPS with tighter batch spacing (225 ns)


Scheduled lift maintenance

Scheduled lift maintenance

  • Some preventive maintenance has been requested on the LHC lifts, and agreed according to the following planning:

    • Tuesday 5 October:             PM15 from 8h to 12h00; PM25 from 13h30 to 17h30

    • Wednesday 6 October:       PX24 from 8h to 12h00; PZ33 from 13h30 to 17h30

    • Thursday 7 October: PM85 from 8h to 12h00; PZ85 from 13h30 to 17h30

    • Tuesday 12 October :          PM56 from 8h to 12h00; PM65 from 13h30 to 17h30

    • Wednesday 13 October:     PZ65 from 8h to 12h00; PM76 from 13h30 to 17h30

    • Thursday 14 October:          PM45 from 8h to 12h00; PZ45 from 13h30 to 17h30

PM56

PZ85

La maintenance du PM56 aura lieu le jeudi 7 de 13h30 à 17h30, celle du PZ 85 le mardi 12 de 8h à 12h.


Access requests

Access Requests

  • TI : demineralized water alarm UA83  Technical stop

  • COD: RCBH15.L6B1

  • RCBXH3.R1 temperature reading issue?


  • Login